# Learning Adaptive and View-Invariant Vision Transformer with Multi-Teacher Knowledge Distillation for Real-Time UAV Tracking

### v1

@@ 3.2 Activation Module (AM)
3.3 View-Invariant Representations (VIR) via Mutual Information Maximization
@@
where $p(\textbf{a},\textbf{b})$ represents the joint probability distribution, while $p(\textbf{a})$ and $p(\textbf{b})$ are the marginal distributions. The symbol $\textup{D}_{KL}$ denotes the Kullback–Leibler divergence (KLD) (MacKay, 2004). Estimating MI is not an easy task in real-world situations, as we usually only have access to the samples that are readily available, not the underlying distributions (Poole et al., 2019). As a result, most existing estimators typically rely on approximating the MI between variables using observed samples. In contrast to these approaches, we employ the Deep InfoMax MI estimator (R.D. and et al., 2019), which estimates MI based on Jensen-Shannon divergence (JSD). This strategy has proven to be effective, as knowing its precise value is less critical than maximizing MI in this context. The Jensen-Shannon MI estimator, represented by $\hat{\textup{I}}_{\mathbf{\mathbf{\Theta}}}^{(JSD)}(\textbf{a},\textbf{b})$, is defined by:
@@
3.4 MI Maximization-based Multi-Teacher Knowledge Distillation (MD)
@@
In the context of multi-teacher knowledge distillation, the selection of teachers-student architectures plays a crucial role in the knowledge distillation process, as it determines how the knowledge from the various teacher models is transferred to the student model. For the teacher models, we use all three off-the-shelf trackers from AVTrack (i.e., AVTrack-DeiT, AVTrack-ViT, and AVTrack-EVA), providing diverse and high-quality teachers without the need for additional operations to implement varied mechanisms in teachers. For the selection of student models, we opt for a self-similar architecture, where the student shares the same structure as the teacher but uses a smaller ViT backbone (i.e., with half ViT blocks). This selection offers two main advantages: 1) it eliminates the need for complex design processes, enabling straightforward implementation and training, and enhancing the interpretability of the model’s behavior; 2) it promotes modularity and scalability, reducing the difficulty of expanding or modifying the model as needed. Averaging the predictions of all teacher models is a common choice (Buciluǎ et al., 2006; Ba and Caruana, 2014; Ilichev et al., 2021; Song and Chai, 2018), as it is a straightforward and effective method. In our implementation, we average the $K$ feature representations $\mathcal{F}_{k}$ produced by the teacher models into an aggregated feature representation $\mathcal{F}=\frac{1}{K}\sum_{k=1}^{K}\mathcal{F}_{k}.$ In practice, to provide richer information during training, the model outputs are softened using a temperature $\tau$, since the original outputs are typically represented in a one-hot encoding format (You et al., 2017). Let $\mathcal{F}^{S}$ represent the feature representation of the student. The softened outputs of the student and teacher are expressed as follows:
@@
where $\tau$ is a constant that we set to 2. Encouraging the student network to replicate the feature representations of the teacher’s final layer network using mean squared error (MSE) loss is straightforward (Ba and Caruana, 2014; Kim et al., 2021; Hamidi et al., 2024). However, since MSE is sensitive to noise and outliers (Hinton et al., 2015), we maximize the MI between the aggregated softened feature of the teacher models and the student model’s softened feature, enhancing generalization and performance, especially in noisy conditions. Given the teachers-student architecture, we apply the Jensen-Shannon MI estimator to achieve MI maximization for multi-teacher knowledge distillation, which establishes the objective function for our MI maximization-based multi-teacher knowledge distillation approach, as expressed in the following equation:
@@
3.5 Prediction Head and Training Objective
Following the corner detection head in (Cui et al., 2022; Ye et al., 2022), we utilize a prediction head $\mathcal{H}$ consisting of several Conv-BN-ReLU layers to directly estimate the target’s bounding box. The output tokens associated with the search image are first reinterpreted into a 2D spatial feature map, which is then input to the prediction head. The head generates three outputs: a target classification score $\mathbf{p}\in[0,1]^{H_{x}/P\times W_{x}/P}$, a local offset $\mathbf{o}\in[0,1]^{2\times H_{x}/P\times W_{x}/P}$, and a normalized bounding box size $\mathbf{s}\in[0,1]^{2\times H_{x}/P\times W_{x}/P}$. The highest classification score is used to estimate the coarse target position, i.e., $(x_{c},y_{c})=\textup{argmax}_{(x,y)}\mathbf{p}(x,y)$, based on which the final target bounding box is determined by:
@@ 3.5 Prediction Head and Training Objective
For the tracking task, we employ the weighted focal loss (Law and Deng, 2018) for classification and a combination of $L_{1}$ loss and GIoU loss (Rezatofighi et al., 2019) for bounding box regression. The overall loss function is formulated as:
@@
where the constants $\lambda_{iou}=2$ and $\lambda_{L_{1}}=5$ are the same as in (Ye et al., 2022), $\gamma$ and $\kappa$ are set to $50$ and $0.0001$, respectively. Our framework is trained end-to-end with the overall loss $\mathcal{L}_{overall}$ after the pretrained weights of the ViT for image classification is loaded. After this training, we apply the proposed MI maximization-based multi-teacher knowledge distillation framework to obtain a student model that better balances performance and efficiency. Specifically, during the distillation phase, the distillation loss $\mathcal{L}_{MD}$ is added to the overall loss from the previous training stage, yielding the total loss $\mathcal{L}_{overall}^{*}$ for end-to-end distillation training as follows:
@@
Table 1: Comparison of precision (Prec.), success rate (Succ.), and speed (FPS) between our trackers and lightweight trackers on DTB70 (Li and Yeung, 2017), UAVDT (Du et al., 2018), VisDrone2018 (Zhu et al., 2018), UAV123 (Mueller et al., 2016), and UAV123@10fps (Mueller et al., 2016). Red , blue , and green signify the first, second, and third places. Please note that the percent symbol (%) is excluded for Prec. and Succ. values.
@@
Method Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. GPU CPU
KCF 46.8 28.0 57.1 29.0 68.5 41.3 52.3 33.1 40.6 26.5 53.1 31.6 - 622.5
fDSST 53.4 35.7 66.6 38.3 69.8 51.0 58.3 40.5 51.6 37.9 60.0 40.7 - 193.4
ECO_HC 63.5 44.8 69.4 41.6 80.8 58.1 71.0 49.6 64.0 46.8 69.7 48.2 - 83.5
MCCT_H 60.4 40.5 66.8 40.2 80.3 56.7 65.9 45.7 59.6 43.4 66.6 45.3 - 63.4
STRCF 64.9 43.7 62.9 41.1 77.8 56.7 68.1 48.1 62.7 45.7 67.3 47.1 - 28.4
ARCF 69.4 47.2 72.0 45.8 79.7 58.4 67.1 46.8 66.6 47.3 71.0 47.1 - 34.2
AutoTrack 71.6 47.8 71.8 45.0 78.8 57.3 68.9 47.2 67.1 47.7 71.6 49.0 - 57.8
DCF-based RACF 72.6 50.5 77.3 49.4 83.4 60.0 70.2 47.7 69.4 48.6 74.6 51.2 - 35.6
HiFT 80.2 59.4 65.2 47.5 71.9 52.6 78.7 59.0 74.9 57.0 74.2 55.1 160.3 -
SiamAPN 78.4 58.5 71.1 51.7 81.5 58.5 76.5 57.5 75.2 56.6 76.7 56.6 194.4 -
P-SiamFC++ 80.3 60.4 80.7 56.6 80.1 58.5 74.5 48.9 73.1 54.9 77.7 55.9 240.5 56.9
TCTrack 81.1 61.9 69.1 50.4 77.6 57.7 77.3 60.4 75.1 58.8 76.0 57.8 139.6 -
TCTrack++ 80.4 61.7 72.5 53.2 80.8 60.3 74.4 58.8 78.2 60.1 77.3 58.8 125.6 -
SGDViT 78.5 60.4 65.7 48.0 72.1 52.1 75.4 57.5 86.3 66.1 75.6 56.8 110.5 -
ABDNet 76.8 59.6 75.5 55.3 75.0 57.2 79.3 60.7 77.3 59.1 76.7 59.1 130.2 -
DRCI 81.4 61.8 84.0 59.0 83.4 60.0 76.7 59.7 73.6 55.2 79.8 59.1 281.3 62.4
CNN-based PRL-Track 79.5 60.6 73.1 53.5 72.6 53.8 79.1 59.3 74.1 58.6 75.2 57.2 135.6 -
Aba-ViTrack 85.9 66.4 83.4 59.9 86.1 65.3 86.4 66.4 85.0 65.5 85.3 64.7 181.5 50.3
LiteTrack 82.5 63.9 81.6 59.3 79.7 61.4 84.2 65.9 83.1 64.0 82.2 62.9 119.7 -
SMAT 81.9 63.8 80.8 58.7 82.5 63.4 81.8 64.6 80.4 63.5 81.5 62.8 124.2 -
LightFC 82.8 64.0 83.4 60.6 82.7 62.8 84.2 65.5 81.3 63.7 82.9 63.4 146.8 -
SuperSBT 84.5 65.4 81.5 60.3 80.4 62.2 85.0 67.2 82.1 65.3 82.7 64.1 121.7 -
@@
AVTrack-DeiT 84.3 65.0 82.1 58.7 86.0 65.3 84.8 66.8 83.2 65.8 84.1 64.4 256.8 59.5
AVTrack-MD-ViT
84.9
65.7
81.4 59.5 84.8 63.7 82.3 65.1
83.5
65.9
83.4 64.0
303.1
63.8
AVTrack-MD-EVA 83.2 63.9 80.8 58.0 84.0 63.5 81.5 62.3 82.7 64.7 82.4 62.5
334.4
67.1
ViT-based AVTrack-MD-DeiT 84.0 65.2
83.1
60.3
84.9
64.2
82.6 65.2 83.3 65.5
83.6
64.1
310.6
64.8
4 Experiments
In this section, we evaluate our tracker on six publicly widely-used UAV tracking benchmarks, including DTB70 (Li and Yeung, 2017), UAVDT (Du et al., 2018), VisDrone2018 (Zhu et al., 2018), UAV123 (Mueller et al., 2016), UAV123@10fps (Mueller et al., 2016), and WebUAV-3M (Zhang et al., 2022a). Specifically, we conduct a thorough comparison with over 38 existing state-of-the-art (SOTA) trackers, using their results obtained from running the official codes with the corresponding hyperparameters. Please note that our evaluation is performed on a PC that was equipped with an i9-10850K processor (3.6GHz), 16GB of RAM, and an NVIDIA TitanX GPU. To provide a clearer evaluation comparison with different types of trackers, we categorize them into two groups: lightweight trackers and deep trackers.
4.1 Implementation Details
@@
Training. For training, we employ a combination of training sets from four different datasets: GOT-10k (Huang et al., 2021), LaSOT (Fan et al., 2019), COCO (Lin et al., 2014), and TrackingNet (Muller et al., 2018). It is noteworthy that all the trackers utilize the same training pipeline to ensure consistency and comparability. The batch size is set to 32. We use the AdamW optimizer to train the model, with a weight decay of $10^{-4}$ and an initial learning rate of $4\times 10^{-5}$. The total number of training epochs is fixed at 300, with 60,000 image pairs processed per epoch. The learning rate decreases by a factor of 10 after 240 epochs. During the multi-teacher distillation phase, we use all three pre-trained AVTrack-* models as teacher models. The parameters of the teacher models are frozen to guide the training of the student model with the proposed knowledge distillation approach, which follows the same training pipeline as that of the teacher models.
Inference. In line with common practices, Hanning window penalties are applied during inference to integrate positional prior on tracking (Zhang et al., 2020). Specifically, a Hanning window of the same size is multiplied by the classification map, and the location with the highest score is selected as the predicted result.
@@
4.2 Comparison with Lightweight Trackers
@@
1) Overall Performance: We compare our methods with 22 SOTA lightweight trackers, inclunding KCF (Henriques et al., 2015), fDSST (Danelljan et al., 2017), ECO_HC (Danelljan et al., 2017), MCCT_H (Wang et al., 2018), STRCF (Li et al., 2018), ARCF (Huang et al., 2019), AutoTrack (Li et al., 2020), RACF (Li et al., 2022b), HiFT (Cao et al., 2021), SiamAPN (Fu et al., 2021), P-SiamFC++ (Wang et al., 2022), TCTrack (Cao et al., 2022), TCTrack++ (Cao et al., 2023), SGDViT (Yao et al., 2023), ABDNet (Zuo et al., 2023), DRCI (Zeng et al., 2023), PRL-Track (Fu et al., 2024), Aba-ViTrack (Li et al., 2023), LiteTrack (Wei et al., 2024), SMAT (Gopal and Amer, 2024), LightFC (Li et al., 2024c), and SuperSBT (Xie et al., 2024a), across five renowned UAV tracking datasets: DTB70 (Li and Yeung, 2017), UAVDT (Du et al., 2018), VisDrone2018 (Zhu et al., 2018), UAV123 (Mueller et al., 2016), and UAV123@10fps (Mueller et al., 2016). The evaluation results are presented in Table 1. From table, our AVTrack-DeiT outperforms all SOTA trackers except Aba-ViTrack (Li et al., 2023) across these five benchmarks in terms of average (Avg.) precision (Prec.) and success rate (Succ.). Specifically, RACF (Li et al., 2022b) exhibits the best performance among DCF-based trackers, with Avg. Prec. and Succ. of 74.6% and 51.2%, respectively. Among CNN-based trackers, DRCI (Zeng et al., 2023) stands out with the highest Avg. Prec. and Succ. of 79.8% and 59.1%. However, even the best methods among DCF- and CNN-based trackers fall short of the worst method among ViT-based trackers. All ViT-based trackers achieve Avg. Prec. and Succ. exceeding 80.0% and 62.0%, respectively, clearly surpassing the CNN-based methods and significantly outperforming the DCF-based ones. Regarding GPU speed, the top three trackers are our proposed methods: AVTrack-MD-EVA, AVTrack-MD-DeiT, and AVTrack-MD-ViT, achieving tracking speeds of 334.4 FPS, 310.6 FPS, and 303.1 FPS, respectively. In terms of CPU speed, all our methods deliver real-time performance on a single CPU [^1] It is important to note that the real-time performance discussed in this work is applicable only to platforms that are similar to or more advanced than ours.. Obviously, DCF-based trackers are the most efficient UAV trackers, as all methods achieving speeds above 80 FPS belong to this category. However, these fastest DCF-based trackers are unable to achieve performance comparable to our methods. For example, the fastest KCF (Henriques et al., 2015) tracker attains only 31.6% in average success rate, which is about half that of our methods. Additionally, DCF-based methods often incur substantial costs to enhance tracking precision. Specifically, RACF exhibits the best performance among DCF-based trackers; however, it runs at only 35.6 FPS, which is significantly lower than the slowest CPU speed of our trackers. Although CNN-based trackers can achieve tracking speeds comparable to our ViT-based trackers, the latter significantly outperform the former in overall performance. Although Aba-ViTrack achieves the highest average precision of 85.3% and the highest average success rate of 64.7%., our AVTrack-DeiT secures second place with only slight gaps of 1.2% and 0.3%, respectively. Furthermore, our AVTrack-DeiT demonstrates higher speeds than Aba-ViTrack. Notably, AVTrack-DeiT runs up to 41%/18% faster than Aba-ViTrack in terms of GPU/CPU speed. When the proposed MI-based multi-teacher knowledge distillation framework (MD) is applied to AVTrack-*, the resulting student model, AVTrack-MD-*, demonstrates a smaller performance drop, with some even achieving performance improvements, all while delivering a significant speed boost. Specifically, AVTrack-MD-* models show varying trade-offs: AVTrack-MD-ViT achieves 0.5% and 0.9% gains in Avg. Prec. and Succ., respectively, with 21.1% GPU and 8.7% CPU speed improvements; AVTrack-MD-EVA gains 0.4% in precision and a minor 0.1% drop in success rate, while also improving GPU and CPU speeds by 17.9% and 6.8%; AVTrack-MD-DeiT has a slight performance drop of 0.5% in precision and 0.3% in success rate but still benefits from significant 20.9% GPU and 8.9% CPU speed gains. These results highlight the advantages of our method and justify its SOTA performance for UAV tracking.
@@
Table 2: Comparison of FLOPs, parameters, inference speed (AGX.FPS) on the NVIDIA Jetson AGX Xavier edge device, and performance on WebUAV-3M (Zhang et al., 2022a).

### v2

@@ III-B Activation Module (AM)
III-C View-Invariant Representations (VIR) via Mutual Information Maximization
@@
where $p(\textbf{a},\textbf{b})$ represents the joint probability distribution, while $p(\textbf{a})$ and $p(\textbf{b})$ are the marginal distributions. The symbol $\textup{D}_{KL}$ denotes the Kullback–Leibler divergence (KLD) [78]. Estimating MI is not an easy task in real-world situations, as we usually only have access to the samples that are readily available, not the underlying distributions [79]. As a result, most existing estimators typically rely on approximating the MI between variables using observed samples. In contrast to these approaches, we employ the Deep InfoMax MI estimator [80], which estimates MI based on Jensen-Shannon divergence (JSD). This strategy has proven to be effective, as knowing its precise value is less critical than maximizing MI in this context. The Jensen-Shannon MI estimator, represented by $\hat{\textup{I}}_{\mathbf{\mathbf{\Theta}}}^{(JSD)}(\textbf{a},\textbf{b})$, is defined by:
@@
III-D MI Maximization-based Multi-Teacher Knowledge Distillation (MD)
@@
In the context of multi-teacher knowledge distillation, the selection of teachers-student architectures plays a crucial role in the knowledge distillation process, as it determines how the knowledge from the various teacher models is transferred to the student model. For the teacher models, we use all three off-the-shelf trackers from AVTrack (i.e., AVTrack-DeiT, AVTrack-ViT, and AVTrack-EVA), providing diverse and high-quality teachers without the need for additional operations to implement varied mechanisms in teachers. For the selection of student models, we opt for a self-similar architecture, where the student shares the same structure as the teacher but uses a smaller ViT backbone (i.e., with half ViT blocks). This selection offers two main advantages: 1) it eliminates the need for complex design processes, enabling straightforward implementation and training, and enhancing the interpretability of the model’s behavior; 2) it promotes modularity and scalability, reducing the difficulty of expanding or modifying the model as needed. Averaging the predictions of all teacher models is a common choice [74, 75, 76, 77], as it is a straightforward and effective method. In our implementation, we average the $K$ feature representations $\mathcal{F}_{k}$ produced by the teacher models into an aggregated feature representation $\mathcal{F}=\frac{1}{K}\sum_{k=1}^{K}\mathcal{F}_{k}.$ In practice, to provide richer information during training, the model outputs are softened using a temperature $\tau$, since the original outputs are typically represented in a one-hot encoding format [68]. Let $\mathcal{F}^{S}$ represent the feature representation of the student. The softened outputs of the student and teacher are expressed as follows:
@@
where $\tau$ is a constant that we set to 2. Encouraging the student network to replicate the feature representations of the teacher’s final layer network using mean squared error (MSE) loss is straightforward [75, 81, 82]. However, since MSE is sensitive to noise and outliers [29], we maximize the MI between the aggregated softened feature of the teacher models and the student model’s softened feature, enhancing generalization and performance, especially in noisy conditions. Given the teachers-student architecture, we apply the Jensen-Shannon MI estimator to achieve MI maximization for multi-teacher knowledge distillation, which establishes the objective function for our MI maximization-based multi-teacher knowledge distillation approach, as expressed in the following equation:
@@
III-E Prediction Head and Training Objective
Following the corner detection head in [14, 12], we utilize a prediction head $\mathcal{H}$ consisting of several Conv-BN-ReLU layers to directly estimate the target’s bounding box. The output tokens associated with the search image are first reinterpreted into a 2D spatial feature map, which is then input to the prediction head. The head generates three outputs: a target classification score $\mathbf{p}\in[0,1]^{H_{x}/P\times W_{x}/P}$, a local offset $\mathbf{o}\in[0,1]^{2\times H_{x}/P\times W_{x}/P}$, and a normalized bounding box size $\mathbf{s}\in[0,1]^{2\times H_{x}/P\times W_{x}/P}$. The highest classification score is used to estimate the coarse target position, i.e., $(x_{c},y_{c})=\textup{argmax}_{(x,y)}\mathbf{p}(x,y)$, based on which the final target bounding box is determined by:
@@ III-E Prediction Head and Training Objective
For the tracking task, we employ the weighted focal loss [83] for classification and a combination of $L_{1}$ loss and GIoU loss [84] for bounding box regression. The overall loss function is formulated as:
@@
where the constants $\lambda_{iou}=2$ and $\lambda_{L_{1}}=5$ are the same as in [12], $\gamma$ and $\kappa$ are set to $50$ and $0.0001$, respectively. Our framework is trained end-to-end with the overall loss $\mathcal{L}_{overall}$ after the pretrained weights of the ViT for image classification is loaded. After this training, we apply the proposed MI maximization-based multi-teacher knowledge distillation framework to obtain a student model that better balances performance and efficiency. Specifically, during the distillation phase, the distillation loss $\mathcal{L}_{MD}$ is added to the overall loss from the previous training stage, yielding the total loss $\mathcal{L}_{overall}^{*}$ for end-to-end distillation training as follows:
@@
TABLE I: Comparison of precision (Prec.), success rate (Succ.), and speed (FPS) between our trackers and lightweight trackers on DTB70 [85], UAVDT [86], VisDrone2018 [87], UAV123 [88], and UAV123@10fps [88]. Red, blue, and green signify the first, second, and third places. Please note that the percent symbol (%) is excluded for Prec. and Succ. values.
@@
Method Source Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. GPU CPU
KCF[89] TAPMI 15 46.8 28.0 57.1 29.0 68.5 41.3 52.3 33.1 40.6 26.5 53.1 31.6 - 622.5
fDSST[90] TPAMI 17 53.4 35.7 66.6 38.3 69.8 51.0 58.3 40.5 51.6 37.9 60.0 40.7 - 193.4
ECO_HC[91] CVPR 17 63.5 44.8 69.4 41.6 80.8 58.1 71.0 49.6 64.0 46.8 69.7 48.2 - 83.5
MCCT_H[92] CVPR 18 60.4 40.5 66.8 40.2 80.3 56.7 65.9 45.7 59.6 43.4 66.6 45.3 - 63.4
STRCF[93] CVPR 18 64.9 43.7 62.9 41.1 77.8 56.7 68.1 48.1 62.7 45.7 67.3 47.1 - 28.4
ARCF[33] ICCV 19 69.4 47.2 72.0 45.8 79.7 58.4 67.1 46.8 66.6 47.3 71.0 47.1 - 34.2
AutoTrack[11] CVPR 20 71.6 47.8 71.8 45.0 78.8 57.3 68.9 47.2 67.1 47.7 71.6 49.0 - 57.8
DCF-based RACF[94] PR 22 72.6 50.5 77.3 49.4 83.4 60.0 70.2 47.7 69.4 48.6 74.6 51.2 - 35.6
HiFT[34] ICCV 21 80.2 59.4 65.2 47.5 71.9 52.6 78.7 59.0 74.9 57.0 74.2 55.1 160.3 -
SiamAPN[95] ICRA 21 78.4 58.5 71.1 51.7 81.5 58.5 76.5 57.5 75.2 56.6 76.7 56.6 194.4 -
P-SiamFC++[36] ICME 22 80.3 60.4 80.7 56.6 80.1 58.5 74.5 48.9 73.1 54.9 77.7 55.9 240.5 56.9
TCTrack[35] CVPR 22 81.1 61.9 69.1 50.4 77.6 57.7 77.3 60.4 75.1 58.8 76.0 57.8 139.6 -
TCTrack++[96] TPAMI 23 80.4 61.7 72.5 53.2 80.8 60.3 74.4 58.8 78.2 60.1 77.3 58.8 125.6 -
SGDViT[97] ICRA 23 78.5 60.4 65.7 48.0 72.1 52.1 75.4 57.5 86.3 66.1 75.6 56.8 110.5 -
ABDNet[98] RAL 23 76.8 59.6 75.5 55.3 75.0 57.2 79.3 60.7 77.3 59.1 76.7 59.1 130.2 -
DRCI[99] ICME 23 81.4 61.8 84.0 59.0 83.4 60.0 76.7 59.7 73.6 55.2 79.8 59.1 281.3 62.4
CNN-based PRL-Track [100] IROS 24 79.5 60.6 73.1 53.5 72.6 53.8 79.1 59.3 74.1 58.6 75.2 57.2 135.6 -
Aba-ViTrack[7] ICCV 23 85.9 66.4 83.4 59.9 86.1 65.3 86.4 66.4 85.0 65.5 85.3 64.7 181.5 50.3
LiteTrack[101] ICRA 24 82.5 63.9 81.6 59.3 79.7 61.4 84.2 65.9 83.1 64.0 82.2 62.9 119.7 -
SMAT[102] WACV 24 81.9 63.8 80.8 58.7 82.5 63.4 81.8 64.6 80.4 63.5 81.5 62.8 124.2 -
LightFC[103] KBS 24 82.8 64.0 83.4 60.6 82.7 62.8 84.2 65.5 81.3 63.7 82.9 63.4 146.8 -
SuperSBT[104] TPAMI 24 84.5 65.4 81.5 60.3 80.4 62.2 85.0 67.2 82.1 65.3 82.7 64.1 121.7 -
@@
AVTrack-DeiT Ours 84.3 65.0 82.1 58.7 86.0 65.3 84.8 66.8 83.2 65.8 84.1 64.4 256.8 59.5
AVTrack-MD-ViT 84.9 65.7 81.4 59.5 84.8 63.7 82.3 65.1 83.5 65.9 83.4 64.0 303.1 63.8
AVTrack-MD-EVA 83.2 63.9 80.8 58.0 84.0 63.5 81.5 62.3 82.7 64.7 82.4 62.5 334.4 67.1
ViT-based AVTrack-MD-DeiT Ours 84.0 65.2 83.1 60.3 84.9 64.2 82.6 65.2 83.3 65.5 83.6 64.1 310.6 64.8
IV Experiments
In this section, we evaluate our tracker on six publicly widely-used UAV tracking benchmarks, including DTB70 [85], UAVDT [86], VisDrone2018 [87], UAV123 [88], UAV123@10fps [88], and WebUAV-3M [105]. Specifically, we conduct a thorough comparison with over 38 existing state-of-the-art (SOTA) trackers, using their results obtained from running the official codes with the corresponding hyperparameters. Please note that our evaluation is performed on a PC that was equipped with an i9-10850K processor (3.6GHz), 16GB of RAM, and an NVIDIA TitanX GPU. To provide a clearer evaluation comparison with different types of trackers, we categorize them into two groups: lightweight trackers and deep trackers.
IV-A Implementation Details
@@
Training. For training, we employ a combination of training sets from four different datasets: GOT-10k [106], LaSOT [107], COCO [108], and TrackingNet [109]. It is noteworthy that all the trackers utilize the same training pipeline to ensure consistency and comparability. The batch size is set to 32. We use the AdamW optimizer to train the model, with a weight decay of $10^{-4}$ and an initial learning rate of $4\times 10^{-5}$. The total number of training epochs is fixed at 300, with 60,000 image pairs processed per epoch. The learning rate decreases by a factor of 10 after 240 epochs. During the multi-teacher distillation phase, we use all three pre-trained AVTrack-* models as teacher models. The parameters of the teacher models are frozen to guide the training of the student model with the proposed knowledge distillation approach, which follows the same training pipeline as that of the teacher models.
Inference. In line with common practices, Hanning window penalties are applied during inference to integrate positional prior on tracking [110]. Specifically, a Hanning window of the same size is multiplied by the classification map, and the location with the highest score is selected as the predicted result.
@@
IV-B Comparison with Lightweight Trackers
@@
1) Overall Performance: We compare our methods with 22 SOTA lightweight trackers, inclunding KCF [89], fDSST [90], ECO_HC [91], MCCT_H [92], STRCF [93], ARCF [33], AutoTrack [11], RACF [94], HiFT [34], SiamAPN [95], P-SiamFC++ [36], TCTrack [35], TCTrack++ [96], SGDViT [97], ABDNet [98], DRCI [99], PRL-Track [100], Aba-ViTrack [7], LiteTrack [101], SMAT [102], LightFC [103], and SuperSBT [104], across five renowned UAV tracking datasets: DTB70 [85], UAVDT [86], VisDrone2018 [87], UAV123 [88], and UAV123@10fps [88]. The evaluation results are presented in Table I. From table, our AVTrack-DeiT outperforms all SOTA trackers except Aba-ViTrack [7] across these five benchmarks in terms of average (Avg.) precision (Prec.) and success rate (Succ.). Specifically, RACF [94] exhibits the best performance among DCF-based trackers, with Avg. Prec. and Succ. of 74.6% and 51.2%, respectively. Among CNN-based trackers, DRCI [99] stands out with the highest Avg. Prec. and Succ. of 79.8% and 59.1%. However, even the best methods among DCF- and CNN-based trackers fall short of the worst method among ViT-based trackers. All ViT-based trackers achieve Avg. Prec. and Succ. exceeding 80.0% and 62.0%, respectively, clearly surpassing the CNN-based methods and significantly outperforming the DCF-based ones. Regarding GPU speed, the top three trackers are our proposed methods: AVTrack-MD-EVA, AVTrack-MD-DeiT, and AVTrack-MD-ViT, achieving tracking speeds of 334.4 FPS, 310.6 FPS, and 303.1 FPS, respectively. In terms of CPU speed, all our methods deliver real-time performance on a single CPU [^1] It is important to note that the real-time performance discussed in this work is applicable only to platforms that are similar to or more advanced than ours.. Obviously, DCF-based trackers are the most efficient UAV trackers, as all methods achieving speeds above 80 FPS belong to this category. However, these fastest DCF-based trackers are unable to achieve performance comparable to our methods. For example, the fastest KCF [89] tracker attains only 31.6% in average success rate, which is about half that of our methods. Additionally, DCF-based methods often incur substantial costs to enhance tracking precision. Specifically, RACF exhibits the best performance among DCF-based trackers; however, it runs at only 35.6 FPS, which is significantly lower than the slowest CPU speed of our trackers. Although CNN-based trackers can achieve tracking speeds comparable to our ViT-based trackers, the latter significantly outperform the former in overall performance. Although Aba-ViTrack achieves the highest average precision of 85.3% and the highest average success rate of 64.7%., our AVTrack-DeiT secures second place with only slight gaps of 1.2% and 0.3%, respectively. Furthermore, our AVTrack-DeiT demonstrates higher speeds than Aba-ViTrack. Notably, AVTrack-DeiT runs up to 41%/18% faster than Aba-ViTrack in terms of GPU/CPU speed. When the proposed MI-based multi-teacher knowledge distillation framework (MD) is applied to AVTrack-*, the resulting student model, AVTrack-MD-*, demonstrates a smaller performance drop, with some even achieving performance improvements, all while delivering a significant speed boost. Specifically, AVTrack-MD-* models show varying trade-offs: AVTrack-MD-ViT achieves 0.5% and 0.9% gains in Avg. Prec. and Succ., respectively, with 21.1% GPU and 8.7% CPU speed improvements; AVTrack-MD-EVA gains 0.4% in precision and a minor 0.1% drop in success rate, while also improving GPU and CPU speeds by 17.9% and 6.8%; AVTrack-MD-DeiT has a slight performance drop of 0.5% in precision and 0.3% in success rate but still benefits from significant 20.9% GPU and 8.9% CPU speed gains. These results highlight the advantages of our method and justify its SOTA performance for UAV tracking.
@@
TABLE II: Comparison of FLOPs, parameters, inference speed (AGX.FPS) on the NVIDIA Jetson AGX Xavier edge device, and performance on WebUAV-3M [105].